OpenStudy (anonymous):

A particle moves along a straight line with equation of motion s = t^{5} - 4 t^{4} Find the value of t (other than 0 ) at which the acceleration is equal to zero.

geerky42 (geerky42):

To find the equation of acceleration, you need to take derivative of equation of motion twice.

geerky42 (geerky42):

Can you find acceleration? @manhani

OpenStudy (anonymous):

20t^3-48t^2

geerky42 (geerky42):

Yeah, looks good. Now set it equals to 0. Can you solve for t?

OpenStudy (anonymous):

do i factor it

geerky42 (geerky42):

Yes. factor \(t^2\) out

OpenStudy (anonymous):

tsq(20t-48)

OpenStudy (anonymous):

you can take out a 4

geerky42 (geerky42):

yeah was about to say that lol

geerky42 (geerky42):

so factor out \(4t^2\)

OpenStudy (anonymous):

4t2(5t-12)

geerky42 (geerky42):

So you have \(4t^2(5t-12)=0\) Either \(4t^2=0\) (t is 0 so we ignore it) or \(5t-12=0\) So solve for t in \(5t-12=0\)

OpenStudy (anonymous):

12/5

OpenStudy (anonymous):

thank you

geerky42 (geerky42):

No problem